OpenStudy (anonymous):

one positive interger is 3 less than another another. the product of the two intergers is 10. what are the intergers? PLEASE HELP !

OpenStudy (anonymous):

let the numbers be x and x-3 x(x-3)=10 x^2-3x-10=0 (x-5)(x+2)=0 x=5 or x=-2 then the numbers are positive so one number=5 other number=(x-3)=(5-3)=2 numbers are 5 & 2